OpenStudy (anonymous):

can someone hlep me solve this by graphing? x-y+3=0 x+y-1=0

OpenStudy (anonymous):

First, solve for y = mx+b, the standard equation for a line. x - y + 3 = 0 -y + 3 = -x -y = -x -3 Multiply by -1 to get rid of the negative y = x + 3 Now solve for y in the next equation x + y -1 = 0 y - 1 = -x y = -x +1 Now, graph these two equations and see where the lines cross. That point is your answer.

OpenStudy (anonymous):

When you graph it, your solution should be (-1,2)

OpenStudy (anonymous):

or x = -1 and y = 2